# Weighted Aggregation of Conformity Scores for Classification

### v1

@@ 2.4 Algorithm for Determining the Weight and the Threshold
3:         Compute score $\langle\bm{w},s(x_{i},y_{i})\rangle$ using $\bm{w}$ and $(x_{i},y_{i})$.
@@
4:              Compute non-conformity score $\langle\bm{w},s(x_{i},y)\rangle$ using $\bm{w}$ and $(x_{i},y)$
@@
For $\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}$ and $\mathcal{I}_{\textrm{test}}$, we present an overview of the data split framework in Figure 1. For what follows, we introduce the following four choices.
@@
    For Validity First Conformal Prediction (VFCP) (Yang & Kuchibhotla, 2024), $\mathcal{I}_{1}=\mathcal{I}_{2}\subseteq\mathcal{I}_{\textrm{train}}$, and $\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}\setminus\mathcal{I}_{1}$. This method divides $\mathcal{I}_{\textrm{train}}$ into two partitions, using less data to determine $\widehat{\bm{w}}$. However, the marginal coverage probability is guaranteed to be at least $1-\alpha$ under the assumption of exchangeability of samples in $\mathcal{I}_{3}$ and $\mathcal{I}_{\textrm{test}}$. The name “validity first” comes from this fact.
@@
    For Efficiency First Conformal Prediction (EFCP) (Yang & Kuchibhotla, 2024), $\mathcal{I}_{1}=\mathcal{I}_{2}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$. This method uses all training data to determine $\widehat{\bm{w}}$, resulting in a more accurate estimation of the optimal $\bm{w}$. However, stronger assumptions are required for the coverage guarantee. The name “efficiency first” comes from this fact.
@@
    For Data Leakage Conformal Prediction (DLCP), $\mathcal{I}_{1}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$, and $\mathcal{I}_{2}=\mathcal{I}_{\textrm{test}}$. Since $\mathcal{I}_{2}$ can be unlabeled data, we can minimize the prediction set on $\mathcal{I}_{\textrm{test}}$ in the step of finding $\widehat{\bm{w}}$ in Eq. (4) by letting $\mathcal{I}_{2}=\mathcal{I}_{\textrm{test}}$. We call this a “data leakage” method because we use the test data in the training procedure of $\widehat{\bm{w}}$.
@@
    For another version of Data Leakage Conformal Prediction (DLCP+), $\mathcal{I}_{1}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{2}=\mathcal{I}_{\textrm{train}}\cup\mathcal{I}_{\textrm{test}}$. If data leakage is allowed, then using the same idea as EFCP, we can use all available data in each step. Therefore, we can include $\mathcal{I}_{\textrm{test}}$ in $\mathcal{I}_{2}$ to maximize the size of samples in the step of finding $\widehat{\bm{w}}$ in Eq. (4).
$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{\textrm{train}}}$$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{3}}$$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{1}}$$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{2}}$ $Q_{1-\alpha}^{(1)}(\bm{w})$ by Eq. (2) $\big{(}\widehat{C}_{1-\alpha}(x_{i},\bm{w},Q_{1-\alpha}^{(1)}(\bm{w}))\big{)}_{i\in\mathcal{I}_{2}}$ by Eq. (3) $\widehat{\bm{w}}$ by Eq. (4) $Q_{1-\alpha}^{(2)}$ by Eq. (5) $\{x_{i}\}_{i\in\mathcal{I}_{\textrm{test}}}$ $\left(\widehat{C}_{1-\alpha}(x_{i},\widehat{\bm{w}},Q_{1-\alpha}^{(2)})\right)_{i\in\mathcal{I}_{\textrm{test}}}$ by Eq. (6)
Figure 1: This example shows an overall framework of data splitting for $\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}$ and $\mathcal{I}_{\textrm{test}}$. The complete procedure is presented in Algorithm 2. Shortly speaking, $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ are used in Step 1 and 2 to select optimal weight $\widehat{\bm{w}}$, $\mathcal{I}_{3}$ is used in Step 3 as calibration set to construct prediction sets for $\mathcal{I}_{\textrm{test}}$. Under this framework, we illustrate four options, VFCP, EFCP, DLCP, and DLCP+, and discuss their coverage and size property in Section 3 as well as empirical results in Section 4.
@@ 3 Theoretical Analysis
The theoretical analysis of our method are based on the subgraph classes in Vapnik–Chervonenkis (VC) theory. This theoretical details and the proofs of lemma and theorems will be left in Section A and B in the Appendix. In this section, we will present the key result for our proposed algorithm. The output of the algorithm analyzed in two aspect:
  1. 1.
    Can the prediction set achieve $1-\alpha$ coverage?
  2. 2.
    Can the prediction set achieve the minimal size? The minimal size will be defined as below.
$\bm{w}^{*}$ is defined as the population level of $\widehat{\bm{w}}.$
$\bm{w}^{*}=\arg\min_{\bm{w}\in\mathcal{W}}\mathbb{E}[|\widehat{C}_{1-\alpha}(X,\bm{w},q)|]\quad{s.t.}\quad\mathbb{P}(Y\in\widehat{C}_{1-\alpha}(X,\bm{w},q))\geq 1-\alpha.$
Equivalently, we can define
$Q_{1-\alpha}(\bm{w})=\sup\{q\in\mathbb{R}:P(Y\in\widehat{C}_{1-\alpha}(X,\bm{w},q))\geq 1-\alpha\},$
@@
$\bm{w}^{*}=\arg\min_{\bm{w}\in\mathcal{W}}\mathbb{E}[|\widehat{C}_{1-\alpha}(X,\bm{w},Q_{1-\alpha}(\bm{w})|]$
Let us define $q^{*}=Q_{1-\alpha}(\bm{w}^{*})$. Since $\bm{w}^{*}$ and $q^{*}$ are deterministic once $\alpha$ and the true distribution are given, We can briefly denote the optimal prediction set for $x$ by
$\widehat{C}^{*}_{1-\alpha}(x):=\widehat{C}_{1-\alpha}(x,\bm{w}^{*},q^{*}).$
From a theoretical perspective, the expected size of the prediction set (at the population level) obtained using $\widehat{\bm{w}}$ cannot be smaller than the one provided by $\bm{w}^{*}$. In the subsequent analysis, we will investigate the difference between the prediction set sizes yielded by $\widehat{\bm{w}}$ and $\bm{w}^{*}$. To facilitate this analysis, we will introduce and rely on the following events. We define
$\Omega(\mathcal{I},\eta)=\sup_{\bm{w}\in\mathbb{R}^{d},q\in\mathbb{R}}\left|\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\mathbf{1}\{\langle\bm{w},s(X_{i},Y_{i})\rangle\geq q\}-\mathbb{E}[\mathbf{1}\{\langle\bm{w},s(X,Y)\rangle\geq q\}]\right|\leq\eta$ (7)
and
$\Gamma(\mathcal{I},\xi)=\sup_{\bm{w}\in\mathbb{R}^{d},q\in\mathbb{R}}\left|\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\sum_{y\in[K]}\mathbf{1}\{\langle\bm{w},s(X_{i},y)\rangle\geq q\}-\mathbb{E}[\mathbf{1}\{\langle\bm{w},s(X,y)\rangle\geq q\}]\right|\leq\xi.$ (8)
$\Omega(\mathcal{I},\eta)$ is about the uniform concentration of the coverage, and $\Gamma(\mathcal{I},\xi)$ is about the uniform concentration of the prediction set size. This is because by the definition of $\widehat{C}_{1-\alpha}$ in Eq. (2) or Eq. (5),
$\mathbf{1}\{\langle\bm{w},s(x,y)\rangle\geq q\}=\mathbf{1}\big{\{}y\in\widehat{C}_{1-\alpha}(x,\bm{w},q)\big{\}}\quad{\text{and}}\quad\sum_{y\in[K]}\mathbf{1}\{\langle\bm{w},s(x,y)\rangle\geq q\}=\big{|}\widehat{C}_{1-\alpha}(x,\bm{w},q)\big{|}.$
@@
$\mathbb{P}\left(\Omega\left(\mathcal{I},8\sqrt{\frac{(d+1)\log(|\mathcal{I}|+1)}{|\mathcal{I}|}}+\delta\right)\right)\geq 1-\exp\left(-\frac{|\mathcal{I}|\delta^{2}}{2}\right).$ (9)
and
$\mathbb{P}\left(\Gamma\left(\mathcal{I},8K\sqrt{\frac{(d+1)\log(|\mathcal{I}|+1)}{|\mathcal{I}|}}+K\delta\right)\right)\geq 1-\exp\left(-\frac{|\mathcal{I}|\delta^{2}}{2}\right).$ (10)
This lemma says that $\Omega(\mathcal{I},\eta)$ and $\Gamma(\mathcal{I},\xi)$ hold with high probability if $\eta,\xi\gtrsim\sqrt{\frac{d\log|\mathcal{I}|}{|\mathcal{I}|}}$. Now we are ready to introduce the statistical guarantees of the methods.
@@ 3.1 Consistency of VFCP
Let $\widehat{C}_{1-\alpha}^{\text{VFCP}}(x)$ be the output of Algorithm 2 with the setting of VFCP, and modified as mention above. Suppose the samples in $\{(X_{i},Y_{i})\}_{i\in\mathcal{I}_{3}\cup\mathcal{I}_{\textrm{test}}}$ are exchangeable, then for $(X,Y)$ in the test set, the coverage probability
$\mathbb{P}\left(Y\in\widehat{C}_{1-\alpha}^{\text{VFCP}}(X)\mid\widehat{\bm{w}}\right)\geq 1-\alpha.$
Moreover, if $\Omega(\mathcal{I}_{1},\eta_{1}),\Omega(\mathcal{I}_{3},\eta_{3})$ and $\Gamma(\mathcal{I}_{2},\xi_{2})$ are satisfied, and $\alpha^{\prime}+\eta_{1}+\eta_{3}\leq\alpha$, then for any $X$ in the test set,
$\mathbb{E}\left[\big{|}\widehat{C}_{1-\alpha}^{\text{VFCP}}(X)\big{|}\mid\widehat{\bm{w}},\widehat{Q}_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})\ \right]\leq\mathbb{E}\left[\big{|}\widehat{C}^{*}_{1-\alpha_{1}}(X)\big{|}\right]+2\xi_{2},$
@@
In the setting of EFCP, $\mathcal{I}_{1}=\mathcal{I}_{2}$. To emphasize their roles in the algorithm, we use different notation in the theorem. $\widehat{\bm{w}}$ and $\widehat{Q}_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})$ depend on the dataset $\mathcal{I}_{\textrm{train}}.$ As long as $\Omega(\mathcal{I}_{1},\eta_{1}),\Omega(\mathcal{I}_{3},\eta_{3})$ and $\Gamma(\mathcal{I}_{2},\xi_{2})$ are satisfied, then the coverage and the expected prediction set size satisfies the inequalities in the theorem.
@@ 3.2 Consistency of EFCP
Let $\widehat{C}_{1-\alpha}^{\text{EFCP}}(x,\widehat{\bm{w}})$ be the output of Algorithm 2 with the setting of EFCP. Suppose the data in $\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{\textrm{test}}$ are independent and $\Omega(\mathcal{I}_{\textrm{train}},\eta_{\text{train}})$ is satisfied. Then for $(X,Y)$ in the test set, the coverage probability
$\mathbb{P}\left(Y\in\widehat{C}_{1-\alpha}^{\text{EFCP}}(X;\widehat{\bm{w}})\mid\widehat{\bm{w}}\ \right)\geq 1-\alpha_{1}\quad\text{where}\quad 1-\alpha_{1}=\frac{1}{|\mathcal{I}_{\textrm{train}}|}\lceil(1+|\mathcal{I}_{\textrm{train}}|)(1-\alpha)\rceil-\eta_{\text{train}}.$
@@
$\mathbb{E}\left[\big{|}\widehat{C}_{1-\alpha}^{\text{EFCP}}(X,\widehat{\bm{w}})\big{|}\mid\widehat{\bm{w}}\ \right]\leq\mathbb{E}\left[\big{|}\widehat{C}^{*}_{1-\alpha_{1}}(X)\big{|}\right]+2\xi_{\text{train}}.$

### v2

@@ 2.4 The Optimal Weight and the Threshold
3:         Compute $\langle\bm{w},s(x_{i},y_{i})\rangle$ using $\bm{w}$ and $(x_{i},y_{i})$.
@@
4:              Compute weighted score $\langle\bm{w},s(x_{i},y)\rangle$
@@
2.5 Data Splitting
In Algorithm 2, the method for splitting the data into $\mathcal{I}_{1}$, $\mathcal{I}_{2}$, and $\mathcal{I}_{3}$ has not been specified. To explore potential data splitting approaches, we first highlight the following two key observations. Firstly, After determining $\widehat{\bm{w}}$, the calibration procedure Eq. (5) and prediction set construction Eq. (6) mirror Step 5 and Step 6 of Algorithm 1. Prior steps aim to identify a weight vector $\bm{w}$ optimizing algorithm performance. Secondly, Eq. (2) and Eq. (5) find quantiles, requiring sample labels. Equations Eq. (3) and Eq. (6) define prediction sets, needing only feature $x$. Thus, the test set can be included in Eq. (3) and Eq. (6).
Based on the observations above, we introduce the following four possible ways of data splitting.
@@ 2.5 Data Splitting
    Validity First Conformal Prediction (VFCP) Yang and Kuchibhotla (2024): $\mathcal{I}_{1}=\mathcal{I}_{2}\subseteq\mathcal{I}_{\textrm{train}}$, and $\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}\setminus\mathcal{I}_{1}$. Divides $\mathcal{I}_{\textrm{train}}$ into two partitions. Marginal coverage probability is guaranteed to be at least $1-\alpha$ under exchangeability of samples in $\mathcal{I}_{3}$ and $\mathcal{I}_{\textrm{test}}$.
@@
    Efficiency First Conformal Prediction (EFCP) Yang and Kuchibhotla (2024): $\mathcal{I}_{1}=\mathcal{I}_{2}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$. Uses all training data to determine $\widehat{\bm{w}}$, resulting in more accurate estimation of optimal $\bm{w}$. This method equires stronger assumptions for coverage guarantee.
@@
    Data Leakage Conformal Prediction (DLCP): $\mathcal{I}_{1}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$, and $\mathcal{I}_{2}=\mathcal{I}_{\textrm{test}}$. Minimizes prediction set on $\mathcal{I}_{\textrm{test}}$ in finding $\widehat{\bm{w}}$ in Eq. (4). Called ”data leakage” as it uses test data in training procedure of $\widehat{\bm{w}}$.
@@
    Data Leakage Conformal Prediction+ (DLCP+): $\mathcal{I}_{1}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{2}=\mathcal{I}_{\textrm{train}}\cup\mathcal{I}_{\textrm{test}}$. Uses all available data in each step, including $\mathcal{I}_{\textrm{test}}$ in $\mathcal{I}_{2}$ to maximize sample size for finding $\widehat{\bm{w}}$ in Eq. (4).
There are various possible methods for data splitting. In the following sections, we focus on these four specific approaches, examining their theoretical properties and evaluating their performance through experiments.
@@ 3 Theoretical Analysis
3.1 Overview of the Results
We investigate the theoretical properties of the proposed methods in terms of validity and efficiency.
Validity: We assess whether the output prediction set achieves the desired coverage rate of $1-\alpha$. For the VFCP method, the coverage rate is guaranteed under the exchangeability assumption. However, for the other three methods, establishing validity is more complex due to the selection bias introduced by $\widehat{\bm{w}}$. Consequently, our aim is to demonstrate that the optimization over $\bm{w}\in\mathcal{W}$ has small impact on validity.
Efficiency: We first observe that the prediction set attains the smallest possible expected size only if the true conditional probabilities $p(y|x)$ are known. Given the predefined score functions, the optimal weight vector in the population sense is defined as
$\bm{w}^{*}=$ $\ \arg\min_{\bm{w}\in\mathcal{W}}\min_{q\in\mathbb{R}}\mathbb{E}\left[|\widehat{C}_{1-\alpha}(X,\bm{w},q)|\right]$
s.t. $\quad\mathbb{P}\left(Y\in\widehat{C}_{1-\alpha}(X,\bm{w},q)\right)\geq 1-\alpha.$
$\bm{w}^{*}$ is the weight vector that minimizes the expected prediction set size while ensuring that, together with an appropriate threshold $q$, the prediction set maintains a desired coverage rate of $1-\alpha$. Equivalently, we can define
$Q_{1-\alpha}(\bm{w}):=\sup\{q\in\mathbb{R}:\mathbb{P}(Y\in\widehat{C}_{1-\alpha}(X,\bm{w},q))\geq 1-\alpha\},$
@@ 3.1 Overview of the Results
$\bm{w}^{*}\in\arg\min_{\bm{w}\in\mathcal{W}}\mathbb{E}\left[|\widehat{C}_{1-\alpha}(X,\bm{w},Q_{1-\alpha}(\bm{w}))|\right].$
We define $q^{*}=Q_{1-\alpha}(\bm{w}^{*})$. Since $\bm{w}^{*}$ and $q^{*}$ are deterministic once $\alpha$ and the true distribution are given, we can succinctly denote the optimal prediction set for $x$:
$\widehat{C}_{1-\alpha}^{*}(x):=\widehat{C}_{1-\alpha}(x,\bm{w}^{*},q^{*}).$
Our analysis of efficiency focuses on the difference between the size of the prediction set produced by the proposed methods using $\widehat{\bm{w}}$ and the expected size of $\widehat{C}^{*}_{1-\alpha}(x)$. In other words, we analyze the discrepancy between the prediction set sizes generated by $\widehat{\bm{w}}$ and the optimal $\bm{w}^{*}$. We will demonstrate that this difference is negligible given a sufficiently large dataset.
We note that this analysis differs from verifying whether $\widehat{\bm{w}}$ converges to $\bm{w}^{*}$, which would require assumptions about the smoothness of the function mapping the weight to the prediction set size. Instead, our analysis makes fewer assumptions on the optimal prediction set size.
3.2 Results by Vapnik–Chervonenkis Theory
We begin by defining the following events on the probability space of pairs of exchangeable random variables $(X_{i},Y_{i})_{i\in\mathcal{I}}$. The event $\Omega(\mathcal{I},\eta)$ is defined as:
$\begin{split}\sup_{\bm{w}\in\mathbb{R}^{d},q\in\mathbb{R}}&\Big{|}\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\mathbf{1}\{\langle\bm{w},s(X_{i},Y_{i})\rangle\geq q\}\\ &-\mathbb{E}[\mathbf{1}\{\langle\bm{w},s(X,Y)\rangle\geq q\}]\Big{|}\leq\eta,\end{split}$ (7)
where $(X,Y)$ has the same distribution as $(X_{1},Y_{1})$. $\Gamma(\mathcal{I},\xi)$ denotes the event
$\begin{split}\sup_{\bm{w}\in\mathbb{R}^{d},q\in\mathbb{R}}\Big{|}&\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\sum_{y\in[K]}\mathbf{1}\{\langle\bm{w},s(X_{i},y)\rangle\geq q\}\\ &-\mathbb{E}[\mathbf{1}\{\langle\bm{w},s(X,y)\rangle\geq q\}]\Big{|}\leq\xi,\end{split}$ (8)
where we recall that $K$ is the number of classes of the classification task. $\Omega(\mathcal{I},\eta)$ is about the uniform concentration of the coverage, and $\Gamma(\mathcal{I},\xi)$ is about the uniform concentration of the prediction set size. This is because by the definition of $\widehat{C}_{1-\alpha}$ in Eq. (2) or Eq. (5),
[…]
$\sum_{y\in[K]}\mathbf{1}\{\langle\bm{w},s(x,y)\rangle\geq q\}=\big{|}\widehat{C}_{1-\alpha}(x,\bm{w},q)\big{|}.$
@@ 3.2 Results by Vapnik–Chervonenkis Theory
  • (a)
    $\Omega\left(\mathcal{I},8\sqrt{\frac{(d+1)\log(|\mathcal{I}|+1)}{|\mathcal{I}|}}+\delta\right)$ hold with probability at least $1-\exp\left(-\frac{|\mathcal{I}|\delta^{2}}{2}\right)$.
  • (b)
    $\Gamma\left(\mathcal{I},8K\sqrt{\frac{(d+1)\log(|\mathcal{I}|+1)}{|\mathcal{I}|}}+K\delta\right)$ hold with probability at least $1-\exp\left(-\frac{|\mathcal{I}|\delta^{2}}{2}\right)$.
The proof of this lemma appears in Section B in the appendix. This lemma establishes that $\Omega(\mathcal{I},\eta)$ and $\Gamma(\mathcal{I},\xi)$ hold with high probability provided that $\eta,\xi\gtrsim\sqrt{\frac{d\log|\mathcal{I}|}{|\mathcal{I}|}}$. The proof, presented in the appendix, applies subgraph classes from Vapnik-Chervonenkis (VC) theory. Additionally, if the inequalities in Equations Eq. (7) and Eq. (8) hold for $\bm{w}\in\mathbb{R}^{d}$, they naturally extend to $\bm{w}\in\mathcal{W}\subset\mathbb{R}^{d}$.
Remark.
Part (b) of Lemma 1 limits the generalization of our current theoretical results to classification tasks. Compared to the bound in Equation Eq. (7), the bound in Equation Eq. (8) involves a summation over $[K]$. In the proof, this is handled using a union bound. However, for regression tasks where prediction set size is typically measured by the Lebesgue measure, a union bound over $K$ is not directly applicable. Therefore, our current work focuses exclusively on classification tasks. Extending the proposed method to regression tasks is left for future research.
@@ 3.3 Consistency of VFCP
Let $\widehat{C}_{1-\alpha}^{\text{VFCP}}(x)$ be the output of Algorithm 2 with the setting of VFCP, and modified as mentioned above. Suppose the samples in $\{(X_{i},Y_{i})\}_{i\in\mathcal{I}_{3}\cup\mathcal{I}_{\textrm{test}}}$ are exchangeable, then for $(X,Y)$ in the test set, the coverage probability
$\mathbb{P}\left(Y\in\widehat{C}_{1-\alpha}^{\text{VFCP}}(X)\right)\geq 1-\alpha.$
Moreover, if $\Omega(\mathcal{I}_{1},\eta_{1}),\Omega(\mathcal{I}_{3},\eta_{3})$ and $\Gamma(\mathcal{I}_{2},\xi_{2})$ are satisfied, and $\alpha^{\prime}+\eta_{1}+\eta_{3}\leq\alpha$, then for test sample $X$,
$\mathbb{E}\left[\big{|}\widehat{C}_{1-\alpha}^{\text{VFCP}}(X)\big{|}\right]\leq\mathbb{E}\left[\big{|}\widehat{C}^{*}_{1-\alpha_{1}}(X)\big{|}\right]+2\xi_{2},$
@@
The proofs of all theorems appear in Section C and Section D and the appendix.
The validity of VFCP requires minimal assumption (exchangeability), while it is less efficient than other method. We will verify this fact in empirical studies.
@@ 3.4 Consistency of EFCP
Let $\widehat{C}_{1-\alpha}^{\text{EFCP}}(x)$ be the output of Algorithm 2 with the setting of EFCP. Suppose the data in $\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{\textrm{test}}$ are independent and $\Omega(\mathcal{I}_{\textrm{train}},\eta_{\text{train}})$ is satisfied. Then for $(X,Y)$ in the test set, the coverage probability
$\mathbb{P}\left(Y\in\widehat{C}_{1-\alpha}^{\text{EFCP}}(X)\right)\geq 1-\alpha_{1}\quad\text{where}$
11 11 $-\alpha_{1}=\frac{1}{|\mathcal{I}_{\textrm{train}}|}\lceil(1+|\mathcal{I}_{\textrm{train}}|)(1-\alpha)\rceil-\eta_{\text{train}}.$
@@
$\mathbb{E}\left[\big{|}\widehat{C}_{1-\alpha}^{\text{EFCP}}(X)\big{|}\right]\leq\mathbb{E}\left[\big{|}\widehat{C}^{*}_{1-\alpha_{1}}(X)\big{|}\right]+2\xi_{\text{train}}.$